OpenStudy (anonymous):

Find the slope of the line that passes through the following points, (-2, 3) and (-1, -3)

OpenStudy (anonymous):

\[gradient/slope=\frac{y2-y1}{x2-x1}\] (-2,3) and (-1,-3) (x1,y1) (x2,y2) now substitute those values into the formula, tell me what you get.

OpenStudy (anonymous):

-6/3 or -2

OpenStudy (anonymous):

\[M = \frac{y2-y1}{x2-x1} = \frac{-3-3}{-1-(-2)}=\frac{-6}{1}=-6\]

OpenStudy (anonymous):

you just need to be careful on the minus part. x2 is -1 and x1 is -2 so when x2 - x1 -1 - (-2) -1 + 2 = 1

hero (hero):

y = mx + b 3 = -2m + b -3 = -1m + b b = 3 + 2m b = m - 3 b = b 3 + 2m = m - 3 2m - m = -3 - 3 m = -6